### v3

# Charting the Landscape of Nefarious Uses of Generative Artificial Intelligence for Online Election Interference

Abstract
Generative Artificial Intelligence (GenAI) and Large Language Models (LLMs) pose significant risks, particularly in the realm of online election interference. This paper explores the nefarious applications of GenAI, highlighting their potential to disrupt democratic processes through deepfakes, botnets, targeted misinformation campaigns, and synthetic identities. By examining recent case studies and public incidents, we illustrate how malicious actors exploit these technologies to try influencing voter behavior, spread disinformation, and undermine public trust in electoral systems. The paper also discusses the societal implications of these threats, emphasizing the urgent need for robust mitigation strategies and international cooperation to safeguard democratic integrity.
Introduction
The advent of GenAI is revolutionizing fields from human-AI teaming to creative content generation, enabling new forms of human-computer interaction and augmenting human capabilities in unprecedented ways (Brown et al., 2020; OpenAI, 2023; Rahwan et al., 2019). However, these advancements come with significant risks: here we focus on the problem of interference with democracy (Bovet and Makse, 2019; Ozmen Garibay et al., 2023) within the broader context of AI ethics (Independent High-Level Expert Group on Artificial Intelligence, 2019; Mittelstadt et al., 2016). The primary objective of this study is to investigate the nefarious applications of GenAI in online election interference, exploring how these technologies can disrupt democratic processes. We aim to provide a comprehensive overview of the threats posed by GenAI, illustrating dual nature paradox that this technology poses, by means of a literature review and analysis to identify key trends and case studies (Barrett et al., 2023).
Online election interference refers to the manipulation of digital information and communication technologies to influence the outcome of elections (Tenove et al., 2018; Kreps and Kriner, 2023). This can include activities such as spreading misinformation, creating fake news, and orchestrating social media campaigns to sway public opinion. The introduction of GenAI into this arena has magnified these threats by enabling more sophisticated and scalable forms of interference (Goldstein et al., 2023; Mozes et al., 2023; Shoaib et al., 2023). Generative AI, with its ability to create realistic text, images, and videos, presents new opportunities for malicious actors to deceive and manipulate the public. For example, deepfake technology can generate realistic videos of political figures saying or doing things they never did, eroding public trust in authentic information sources (see Fig. 1) (Seymour et al., 2023; Barrett et al., 2023; Kenthapadi et al., 2023; Lorenz et al., 2023; Gabriel et al., 2024). Moreover, the widespread use of AI-powered botnets can amplify misinformation and create the illusion of widespread support or opposition to certain political ideas, further distorting the democratic process (Unver, 2018; Yang and Menczer, 2023; Kreps and Kriner, 2023; Coeckelbergh, 2023).
In our review, we identified several key case studies that illustrate the impact of GenAI on election interference (Yang et al., 2019; Kreps et al., 2022; Chennupati, 2024). For instance, during the 2016 U.S. presidential election, AI-generated botnets played a significant role in spreading misinformation and influencing public opinion (Bessi and Ferrara, 2016; Howard et al., 2018; Shao et al., 2018; Boichak et al., 2018). Similarly, deepfake videos have been used to create false narratives about political candidates, affecting voter perceptions and trust (Vaccari and Chadwick, 2020; Dobber et al., 2021; Diakopoulos and Johnson, 2021; Seymour et al., 2023). The deployment of GenAI in online election interference poses a unique challenge to democracies worldwide. Traditional methods of combating misinformation and digital manipulation are often insufficient to address the sophisticated techniques enabled by GenAI. This necessitates a comprehensive understanding of the various ways in which GenAI can be exploited for nefarious purposes and the development of new strategies to mitigate these risks (Xu et al., 2023; Gupta et al., 2023; Ziems et al., 2023).
This perspective article aims to shed light on the darker applications of GenAI in the context of online election interference. By examining the various nefarious uses of these technologies, we hope to highlight the urgent need for regulatory oversight, technological solutions, public awareness, and collaborative efforts to safeguard the integrity of democratic processes. Through this exploration, we seek to provide a foundation for policymakers, technologists, and civil society to develop robust strategies to protect elections from AI-driven interference, ensuring that the benefits of GenAI are not overshadowed by its potential for harm (Rudin, 2019; Cao et al., 2023).
The Dual Nature of Generative AI
GenAI and LLMs, celebrated for their ability to process natural language and generate multimodal content, offer numerous benefits, including enhancing creativity and solving complex problems (Brown et al., 2020; OpenAI, 2023; Shen et al., 2024; Yao et al., 2024). These technologies have enabled advancements in various fields, from natural language understanding and translation to creative industries such as art and music (Cao et al., 2023; Franceschelli and Musolesi, 2023). However, these same capabilities can be exploited for malicious purposes, such as creating deepfakes, synthetic identities, and targeted misinformation campaigns (Barrett et al., 2023; Mozes et al., 2023; Kang et al., 2024).
On the positive side, GenAI can assist in drafting documents, writing code, generating realistic images and videos, and even conducting complex data analysis (Mandapuram et al., 2018; Bail, 2024; Li et al., 2024). These applications have immense potential to boost productivity and innovation across different sectors (Ooi et al., 2023; Kanbach et al., 2024). For example, AI-generated content can support marketing efforts by creating personalized advertisements or generating engaging social media posts. In healthcare, GenAI can help in diagnosing diseases by analyzing medical images or generating detailed patient reports (Davenport and Kalakota, 2019; Sai et al., 2024).
Despite these benefits, the dark side of GenAI cannot be ignored (Wach et al., 2023; Kang et al., 2024). The same technology that generates realistic images and videos for legitimate purposes can also create deepfakes, which are manipulated media that convincingly alter appearances or actions of individuals. Deepfakes can be used to spread false information, damage reputations, and even blackmail individuals (Vaccari and Chadwick, 2020; Diakopoulos and Johnson, 2021; Seymour et al., 2023). The ease with which these deepfakes can be created poses a significant threat to the authenticity of information shared online. Similarly, LLMs can be used to generate convincing fake news articles, social media posts, and even entire personas (Chen and Shu, 2023; Augenstein et al., 2024; Choi and Ferrara, 2024a, b; Quelle and Bovet, 2024). These synthetic identities can be employed to infiltrate social networks, manipulate public opinion, and disrupt democratic processes (Haider et al., 2023; Luceri et al., 2024). AI-generated text can mimic human writing so effectively that distinguishing between genuine and fake content becomes challenging (Zellers et al., 2019; Augenstein et al., 2024; Quelle and Bovet, 2024; Jones and Bergen, 2024). Moreover, GenAI’s ability to tailor messages to specific demographics makes it a powerful tool for targeted misinformation campaigns. These campaigns can be designed to exploit societal divisions, amplify existing biases, and influence voter behavior (Boichak et al., 2018; Sap et al., 2019; Bovet and Makse, 2019; Goldstein et al., 2023; Ezzeddine et al., 2023; Hofmann et al., 2024). The targeted nature of these campaigns increases their effectiveness and the potential harm they can cause.
The dual nature of GenAI highlights the need for a balanced approach in leveraging these technologies. While the benefits are substantial, the risks are equally significant (Weidinger et al., 2021, 2022; Barrett et al., 2023; Ferrara, 2024b). As we continue to develop and deploy GenAI, it is crucial to implement safeguards that prevent its misuse (Rudin, 2019). This includes developing robust detection mechanisms, establishing ethical guidelines, and promoting public awareness about the potential dangers of AI-generated content (Mittelstadt et al., 2016; Shahriari and Shahriari, 2017; Independent High-Level Expert Group on Artificial Intelligence, 2019; Jobin et al., 2019; Smuha, 2019; Yang et al., 2019; Qiu et al., 2023; Zhou et al., 2023; Sison et al., 2023). In conclusion, while GenAI holds great promise for enhancing various aspects of society, it also presents substantial risks. Understanding the dual nature of these technologies is essential for developing strategies that maximize their benefits while mitigating their potential harms (Seymour et al., 2023; Mozes et al., 2023).
Survey Approach
To sift through the growing scholarly work about the risks of GenAI, we carried out a systematic literature review and analysis. The literature review involved systematically identifying and analyzing scholarly articles, reports, and case studies related to the use of GenAI in online election interference. We focused on sources that provide empirical evidence and theoretical insights into the nefarious applications of these technologies. The analysis was conducted to identify key trends, influential works, and research gaps in the field. We analyzed recent journal and conference publication patterns, for a comprehensive understanding of the topic’s evolution.
Table 1. Summary of Nefarious Applications of GenAI in Online Election Interference
Nefarious Application Mechanism Impact Examples Countermeasures Future Trends Deepfakes AI-generated realistic videos/audio Erosion of public trust, reputation damage Fake celebrity endorsements, political smear campaigns AI-based detection tools, legal regulations More sophisticated, harder-to-detect deepfakes AI-Powered Botnets Automated social media accounts Distortion of public opinion, misinformation spread Bot campaigns during elections, coordinated disinformation Bot detection algorithms, platform policies Increasingly human-like behavior, decentralized bot networks Targeted Misinformation Campaigns Tailored false content Exploitation of societal divisions, voter manipulation False news articles, doctored images Fact-checking services, user education AI-generated personalized misinformation Synthetic Identities AI-created personas Infiltration of online communities, intelligence gathering Fake profiles on social media, false expert identities Identity verification protocols, AI detection tools More convincing and diverse synthetic identities
Nefarious Applications in Election Interference
The integration of GenAI into the sphere of online election interference introduces numerous sophisticated threats to the integrity of democratic processes. These applications leverage the advanced capabilities of AI to manipulate information and influence voter behavior in unprecedented ways. Here, we explore the primary nefarious applications of GenAI in election interference (Table 1).
Deepfakes and Synthetic Media
One of the most alarming applications of GenAI is the creation of deepfakes and other forms of synthetic media. Deepfakes use AI to create hyper-realistic videos and audio recordings of individuals, often portraying them saying or doing things they never did. Fabricated content can be used to spread false information, damage reputations, and manipulate public opinion (Seymour et al., 2023).
The potential impact of deepfakes on elections is significant. For instance, a deepfake video of a political candidate making controversial statements or engaging in unethical behavior could go viral, swaying public perception and potentially altering the outcome of an election. The speed at which such content can spread on social media platforms exacerbates the problem, making it difficult to counteract the false information before it influences a large audience (Vosoughi et al., 2018; Pinto et al., 2024).
Botnets and Social Media Manipulation
AI-powered botnets represent another potent tool for election interference. Botnets are networks of automated accounts that can post, like, share, and comment on social media platforms, creating the illusion of widespread support or opposition for certain viewpoints (Yang and Menczer, 2023). These bots can be programmed to spread misinformation, amplify divisive content, and suppress legitimate discourse, thereby distorting the digital public sphere (Ferrara et al., 2016).
The sophistication of modern AI allows these bots to mimic human behavior convincingly, making it challenging to detect and eliminate them (Ferrara, 2023b). For example, during the 2016 U.S. presidential election, bots were used extensively to influence online discussions and spread false information, contributing to the overall misinformation ecosystem (Ferrara et al., 2016; Ferrara, 2018). Pozzana and Ferrara (2020) provided insights into the dynamics between bot and human behaviors, illustrating the increasing sophistication of bot activities; Luceri et al. (2019b, a) showed how these behaviors evolve as a function of political ideology. The use of such technology in future elections poses a continuing threat to the integrity of democratic processes (Yang et al., 2022).
Targeted Misinformation Campaigns
GenAI’s ability to synthesize realistic and persuasive text enables the creation of highly effective targeted misinformation campaigns. These campaigns are designed to exploit existing societal divisions and biases, tailoring messages to specific demographic groups to maximize their impact (Vosoughi et al., 2018). By leveraging data from social media platforms and other online sources, malicious actors can create personalized propaganda that resonates deeply with targeted individuals, influencing their opinions, emotions, and behaviors.
For example, during an election, a targeted misinformation campaign might disseminate false information about a candidate’s policies or personal life, aiming to dissuade certain voter groups from supporting them. The personalized nature of these messages makes them particularly effective, as they can address the specific concerns and biases of the recipients (Allcott and Gentzkow, 2017). This form of manipulation undermines the democratic process by skewing the electorate’s perception based on falsehoods.
Synthetic Identities and Fake Accounts
GenAI can also be used to create convincing synthetic identities and fake accounts, which can be employed to infiltrate online communities, spread misinformation, and gather intelligence on political opponents (Ferrara, 2023b, 2024a). These synthetic identities can come with complete profiles, including photos, bios, and posting histories, making them difficult to distinguish from real users (Westerlund, 2019; Ezzeddine et al., 2023). The creation of fake accounts can facilitate a range of malicious activities. For instance, these accounts can be used to spread false narratives, engage in coordinated attacks against political figures, or even influence online polls. The use of synthetic identities to manipulate online discourse presents a significant challenge for social media platforms and requires sophisticated detection and countermeasures (Haider et al., 2023).
In summary, the nefarious applications of GenAI in election interference are diverse and highly sophisticated. From creating deepfakes to orchestrating botnets, these technologies offer malicious actors powerful tools to disrupt democratic processes and influence voter behavior. Addressing these threats requires a multifaceted approach that includes technological solutions, regulatory measures, and public awareness initiatives (Ferrara, 2024a).
Table 2. Examples of Past Election Interference Operations Enabled by AI or Bots.
Election Country Techniques Used Actors Involved Outcome References 2016 UK Bots, social media campaigns, fake news Pro-Brexit groups, foreign actors Influenced public opinion, supported Brexit campaign Howard and Kollanyi (2016) 2016 USA Botnets, fake news, social media manipulation Russian state-sponsored groups Influenced public opinion, increased polarization Ferrara et al. (2016); Vosoughi et al. (2018) 2016 Philippines Troll accounts, fake news production Domestic political actors, misinformation agents Manipulated public opinion, supported political candidates Ong and Cabanes (2018); Haider et al. (2023) 2017 France Deepfakes, hacked emails, social media campaigns Russian state-sponsored groups, domestic actors Attempted to discredit candidates, disrupted public discourse Ferrara (2017); Pierri et al. (2023) 2017 Spain Bots, fake news, social media manipulation Pro-separatist groups, foreign actors Influenced public opinion, supported Catalan independence Stella et al. (2018) 2017 Germany Bots, trolls, misinformation Russian state-sponsored groups Influenced public opinion, attempted to disrupt political stability Bennhold and Barry (2017) 2018 USA Social media bots, fake news, targeted ads Foreign actors, domestic political groups Influenced voter behavior, increased polarization Deb et al. (2019) 2014/16 & 2018/22 Brazil Social media bots, fake news, WhatsApp campaigns Domestic political groups Polarized electorate, influenced election results Arnaudo (2017); Pacheco (2023) 2018 Italy Social media bots, misinformation campaigns Various political actors, misinformation agents Influenced public opinion, affected election outcomes Caldarelli et al. (2020) 2019 Indonesia Spear phishing, cyber attacks, social media manipulation Various political actors, cyber agents Influenced voter behavior, disrupted electoral process Tapsell (2020) 2019 India and Pakistan Fake social media accounts, coordinated misinformation Various political parties, independent actors Spread misinformation, manipulated voter sentiment Anand and Jaggi (2019) 2020 USA Deepfakes, AI-generated fake news, targeted ads Foreign actors, political groups Increased misinformation, voter suppression efforts Ferrara (2020); Ferrara et al. (2020) 2020 Taiwan Digital civic participation, misinformation campaigns Various political actors, misinformation agents Spread of misinformation, influence on voter perception and behavior Chang et al. (2021)
The Societal Implications
The misuse of GenAI in election interference poses several significant societal risks. These risks extend beyond the immediate impact on elections (cf. Tables 2 and 3) and their outcomes and have profound implications for democratic institutions, social cohesion, and public trust.
Erosion of Public Trust
The widespread use of deepfakes and misinformation can severely erode public trust in media, institutions, and the democratic process itself. When voters are exposed to manipulated content that appears authentic, their ability to discern truth from falsehood is compromised. This erosion of trust can lead to cynicism and apathy among the electorate, as people become increasingly skeptical of all information sources (Seymour et al., 2023). The impact of this distrust is far-reaching. It undermines the credibility of legitimate news organizations and government institutions, making it easier for malicious actors to spread further misinformation (Baribi-Bartov et al., 2024). Once trust is broken, restoring it is complex, requiring significant efforts from involved stakeholders like media organizations, governments, and broader society (Ferrara, 2024a).
Polarization and Division
Targeted misinformation campaigns exacerbate existing societal divisions by exploiting and amplifying biases and prejudices. By delivering tailored messages that resonate with specific demographic groups, these campaigns can deepen political and social polarization. For instance, misinformation targeting different ethnic or social groups can inflame tensions and lead to increased hostility between communities (Vosoughi et al., 2018). Such polarization weakens the social fabric, making it more challenging to achieve consensus on critical issues. This division can manifest in various forms, from increased partisanship in political discourse to violent confrontations between opposing groups. The long-term consequences of such division are detrimental to social stability and cohesion (Allcott and Gentzkow, 2017).
Undermining Democracy
The ability to manipulate public opinion on a large scale poses a direct threat to the foundations of democracy. Elections are meant to reflect the informed choices of the electorate. However, when these choices are influenced by false information, the legitimacy of the electoral process is compromised. This manipulation undermines the principle of free and fair elections, a cornerstone of democratic governance (Vosoughi et al., 2018). Badawy et al. (2019b) examined the demographics of individuals most susceptible to online political manipulation, highlighting the vulnerability of certain groups.
Furthermore, the perception that elections can be easily manipulated through AI-driven misinformation campaigns can lead to decreased voter turnout. If people believe their votes do not matter or that the process is rigged, they are less likely to participate in the democratic process. This disengagement weakens democratic institutions and can lead to governance that is less representative of the people’s will (Ferrara, 2023b).
Exacerbation of Inequality
The deployment of GenAI technologies in ways that disproportionately affect certain groups can exacerbate existing social inequalities (Ferrara, 2023a). For example, misinformation campaigns might target marginalized communities with content designed to suppress their voter turnout or sow confusion about voting processes (Calo, 2017). Such tactics exploit the vulnerabilities of these communities, further entrenching their disenfranchisement. Moreover, access to sophisticated AI technologies is often concentrated among well-resourced entities, such as state actors or large corporations, giving them disproportionate power to influence public opinion and elections. This concentration of power can lead to a democratic deficit, where the interests of a few overshadow the voices of the many (Gillespie, 2018).
Psychological Impact on Society
The constant exposure to AI-generated misinformation and deepfakes can have a psychological toll on the populace. The bombardment of conflicting information can lead to information overload, where individuals struggle to process and make sense of the vast amounts of data they encounter daily: This situation can cause increased stress, anxiety, and a sense of helplessness among individuals (Fu et al., 2020; Matthes et al., 2020; Brooks et al., 2024). Additionally, the fear of being deceived by AI-generated content can lead to a general mistrust of digital information. This mistrust can hinder the adoption of beneficial technologies and stifle innovation, as people become wary of engaging with digital platforms (Ferrara, 2024a).
Table 3. Examples of Research on Nefarious Applications of AI for Election Interference
Study Main Findings
Ratkiewicz et al. (2011) Introduces the "Truthy" project, which detects and tracks political smears, astroturfing, misinformation, and other political abuse in social media.
Bessi and Ferrara (2016) Found that social bots – later associated with Russia’s IRA (Broniatowski et al., 2018; Badawy et al., 2019a) – were used to significantly influence the 2016 U.S. Presidential election discussions on Twitter.
Shao et al. (2018) Identified that social bots amplify low-credibility content and contributed to the spread of misinformation during the 2016 U.S. election.
Im et al. (2020) Demonstrated that Russian troll accounts continue to operate and adapt their strategies post-2016 U.S. elections.
Haider et al. (2023) Developed a framework to demonstrate an approach to detect online political abuse in low-resource languages.
Stella et al. (2018) Showed that bots disproportionately amplify negative and inflammatory content, leading to greater exposure and potential influence on public opinion.
Chang et al. (2021) Explored how digital civic participation and misinformation campaigns affected the 2020 Taiwanese Presidential election.
Luceri et al. (2019b) Analyzed the behavior of bots and humans on social media during elections, finding that bots become more sophisticated and difficult to detect over time.
Mitigation Strategies
To combat the multifaceted threats posed by GenAI in online election interference, a comprehensive and multi-pronged approach is essential. This section outlines various strategies, including regulatory measures, technological solutions, public awareness campaigns, and collaborative efforts, aimed at mitigating the risks associated with the misuse of GenAI.
Regulation and Oversight
Effective regulation and oversight are crucial in governing the use of GenAI, particularly in sensitive areas like elections. Governments and international bodies need to establish clear guidelines and standards for the ethical use of AI technologies. These regulations should include provisions for transparency, accountability, and the prevention of misuse (Floridi, 2019).
One approach is the implementation of AI ethics guidelines, such as the European Union’s Ethics Guidelines for Trustworthy AI, which emphasize principles like human agency, technical robustness, privacy, and accountability (Independent High-Level Expert Group on Artificial Intelligence, 2019). Additionally, regulatory frameworks should mandate the disclosure of AI-generated content to ensure that the public can identify and scrutinize such content appropriately (OpenAI, 2019).
Technological Solutions
Advances in technology can provide robust solutions to detect and mitigate the impact of AI-generated misinformation. For instance, digital watermarking and forensic techniques can be employed to identify deepfakes and other synthetic media (Perov et al., 2020). These technologies can help verify the authenticity of digital content, making it easier to distinguish genuine from manipulated media.
Moreover, AI-driven detection systems can be developed to identify and neutralize botnets and fake accounts on social media platforms. These systems can use machine learning algorithms to detect patterns indicative of automated behavior, enabling platform operators to take swift action against malicious bots (Yang et al., 2022). Implementing such technologies requires collaboration between tech companies, researchers, and policymakers to ensure their effectiveness and scalability.
Public Awareness and Education
Educating the public about the potential for AI-driven misinformation and how to critically evaluate information sources is crucial in combating the spread of false information. Public awareness campaigns can help individuals develop the skills needed to identify deepfakes, recognize signs of bot activity, and verify the credibility of news sources (Rudin, 2019).
Educational initiatives should be integrated into school curricula and adult education programs to promote media literacy. These programs can teach critical thinking skills and provide tools for assessing the reliability of information online (Ferrara, 2023b). Additionally, public service announcements and partnerships with media organizations can amplify these efforts, reaching a broader audience.
Collaborative Efforts
Addressing the challenges posed by GenAI requires collaborative efforts among policymakers, technologists, and civil society. Multi-stakeholder initiatives can foster the development of best practices and shared standards for the ethical use of AI. For example, the Partnership on AI brings together diverse organizations to advance public understanding of AI and promote the responsible use of AI technologies (Heer, 2018).
Moreover, international cooperation is essential to address the global nature of online election interference. Governments and international organizations should work together to share information, coordinate responses, and develop joint strategies to counteract AI-driven threats. This collaboration can help build a unified front against the misuse of GenAI, ensuring that efforts are aligned and resources are effectively utilized (Taddeo and Floridi, 2018).
Ethical AI Development
Promoting ethical AI development involves ensuring that AI systems are designed with fairness, accountability, and transparency in mind. Developers and researchers should adhere to ethical guidelines and consider the societal impact of their work. This includes conducting thorough risk assessments, implementing bias mitigation strategies, and fostering inclusive design practices (Weidinger et al., 2022).
Ethical AI development also involves creating AI systems that are interpretable and explainable. Interpretable models allow users to understand how decisions are made, increasing trust in AI systems and enabling more effective oversight (Rudin, 2019). By prioritizing ethical considerations throughout the AI development lifecycle, developers can help prevent the misuse of AI technologies in election interference and other contexts (Bellamy et al., 2019).
Legal and Policy Frameworks
Developing robust legal and policy frameworks is essential to address the challenges posed by GenAI. These frameworks should encompass data protection laws, cybersecurity measures, and regulations specific to AI technologies. For instance, the General Data Protection Regulation (GDPR) in the European Union provides a strong foundation for data privacy and security, which is crucial in preventing unauthorized access and manipulation of data used in AI systems (Goodman and Flaxman, 2017).
Policies should also address the accountability of AI developers and users, ensuring that there are consequences for the misuse of AI technologies. This can include penalties for creating and disseminating deepfakes, as well as requirements for transparency in AI-generated content. Legal frameworks need to be adaptive and responsive to the rapidly evolving AI landscape, ensuring they remain effective in mitigating emerging threats (Mittelstadt et al., 2016).
In conclusion, addressing the societal implications of GenAI in election interference requires a comprehensive approach that integrates regulatory measures, technological solutions, public education, collaborative efforts, ethical development practices, and robust legal frameworks. By implementing these strategies, we can mitigate the risks associated with GenAI and protect the integrity of democratic processes (Ferrara, 2024b).
Conclusions
This study highlights the significant risks posed by Generative Artificial Intelligence (GenAI) in the context of online election interference. By employing a dual methodology of literature review and analysis, we have provided a comprehensive overview of the nefarious applications of GenAI. Our findings underscore the urgent need for robust mitigation strategies, including regulatory measures, technological solutions, public awareness campaigns, and international cooperation.
Our research contributes to the existing body of knowledge by elucidating the specific threats posed by GenAI and proposing actionable recommendations to safeguard democratic processes. Future research should focus on developing advanced detection mechanisms and exploring the ethical implications of GenAI to ensure its responsible use.
Data Availability
Data sharing is not applicable.
About the Team
The 2024 Election Integrity Initiative is led by Emilio Ferrara and Luca Luceri and carried out by a collective of USC students and volunteers whose contributions are instrumental to enable these studies. The author is grateful to the following HUMANS Lab’s members for their tireless efforts on this project: Ashwin Balasubramanian, Leonardo Blas, Charles ’Duke’ Bickham, Keith Burghardt, Sneha Chawan, Vishal Reddy Chintham, Eun Cheol Choi, Srilatha Dama, Priyanka Dey, Isabel Epistelomogi, Saborni Kundu, Grace Li, Richard Peng, Gabriela Pinto, Jinhu Qi, Ameen Qureshi, Namratha Sairam, Tanishq Salkar, Srivarshan Selvaraj, Kashish Atit Shah, Gokulraj Varatharajan, Reuben Varghese, Siyi Zhou, and Vito Zou.
Further Readings